# Towards Scalable and Stable Parallelization of Nonlinear RNNs

### v1

@@ Appendix B Experimental Details
B.1 Quasi-DEER for Training
In this section we elaborate on our Experiment 1, discussed in Section 6.1. We hew closely to the experimental design of Section 4.1 of Lim et al. [2024], including 5 warm-up steps for all timing experiments. However, instead of running 5 seeds for 20 repetitions each, we run 20 seeds for 5 repetitions each, to get more converage over different evaluation (as the timing for each evaluation show low variance). We also include memory profiling experiments not present in Lim et al. [2024], which we discuss in more detail in Section B.1.3.
@@ B.1.1 Numerical Precision of DEER and Quasi-DEER
Figure 5: The accuracy of evaluating with parallelized methods (DEER and quasi-DEER) as opposed to sequential evaluation. The parallelized methods to the correct trace within numerical precision. The hidden state size is $D=4$ and the sequence length is $T=10,000$.
@@ B.1.2 Different Scaling Regimes Depending on GPU Saturation
In Figure 6, we run the timing benchmarks of Section 6.1 on a wider range of sequence lengths $T$ and hidden state sizes $D$, on a large GPU (a V100 with 32 GB) and a smaller batch size (of 1). In doing so, we highlight that the parallel nature of DEER and quasi-DEER, as their wall-clock time scales sublinearly in the sequence length $T$ in smaller ($D$, $T$) regimes. However, we note that in the larger regimes considered in our main text and in Lim et al. [2024], we often observe linear scaling in the sequence length $T$ for the wall-clock time of DEER and quasi-DEER, even though these algorithms are still faster than sequential evaluation. Figure 6 shows good evidence that these parallel algorithms are suffering from saturation of the GPU, and would benefit from even more optimized parallel hardware.
Figure 6: Evaluating an untrained GRU. Sublinear and linear timing regimes for parallelized algorithms. The above experiments were run on a 32 GB V100 with a batch size of 1. As in Figure 2, we use 20 seeds for timing, 3 sees for memory, and the dashed black line indicates the memory capacity of the GPU (32 GB). We observe that in smaller regimes in $D$ and $T$ that the wall-clock time shows sublinear scaling indicative of the use of parallel algorithms. However, when the GPU rank becomes saturated, the benefits of parallelization are reduced and we begin to see linear scaling in wall-clock time with $T$.
@@ B.1.3 Memory Profiling Details
B.2 Quasi-DEER for Evaluation
Here we discuss the experimental details for the experiment in Section 6.2. We follow the same experimental set up as in Section 4.3 and Appendix B.3 of Lim et al. [2024]. As an aside, we note that the choice of hardware can impact behavior of the algorithms dramatically. For replicability, we run on the same hardware as Lim et al. [2024], using a 16GB V100 SXM2. However, we note that if we try to run these same experiments on A100, DEER struggles to converge numerically, although quasi-DEER shows no such difficulty. If we run on a CPU, both DEER and quasi-DEER converge numerically. On balance, on the eigenworms time series classification task, both DEER and quasi-DEER are numerically stable for the most part; the numerical instabilities we have observed for DEER on an A100 are likely specific to some particular detail of JAX/hardware interaction.
@@ B.3.1 AR GRU Architecture
[…]
This AR GRU is trained using standard sequential evaluation and backpropagation-through-time to produce a noisy sine wave of length $10,\!000$. We train the AR GRU on $1024$ traces $x_{1:T}$ generated from a sine wave with amplitude 10 and white noise applied to each time step, and the training objective is to minimize the the negative log probability of the $x_{1:T}$.
Once the AR GRU has been trained, it can generate its own trace $\tilde{x}_{1:T}$ given an initial hidden state $h_{0}$ and noises $\epsilon_{1:T}$.
We note that such a system is Markovian with dimension $D=\dim(h)+\dim(x)$, as together the hidden state $h_{t}$ and output $x_{t+1}$ determine the next hidden state $h_{t+1}$ and output $x_{t+2}$. Thus, in the notation of Section 2, a hidden state $\mathbf{s}_{t}$ of the Markovian state space model is $\mathbf{s}_{t}=(x_{t+1},h_{t})$. Therefore, we can apply fixed point methods to try to find the correct trace $\mathbf{s}^{*}$ in a parallelized manner instead of autoregressively.
@@ B.3.2 Wall-clock Time Benchmark
The timing experiments were carried out as follows on an A100 GPU. We ran sequential evaluation of the trained AR GRU to produce noisy sine waves of length $T=10,\!000$, as well as the four parallelized method we consider in this paper (DEER, quasi-DEER, ELK, and quasi-ELK).
We ran 20 different random seeds (which lead to different values of the $\epsilon_{1:T}$ and therefore different nonlinear dynamics), and timed each for a total of 4 repetitions. Thus, for each of the five methods, we have a total of 80 timing runs. We record the wall-clock time needed to evaluate the length $T$ sequence sequentially, as well as wall-clock time, divided by $T$ needed to run $T$ Newton iterations of each of the parallelized methods (thus, we obtain the time per newton iteration for each of the parallelized methods).
Over these 80 timing runs, the sequential evaluation took an average of 96.2 milliseconds, with standard deviation of $1.55$ ms. We report the average time per newton iteration, the total number of newton iterations needed for convergence, and the total wall-clock time to convergence in Table 2. Note that the third column of Table 2 is the product of the first two columns.
We effectively read the number of Newton iteration to convergence off of the graphs in Figure 4, but find the number of Newton iterations to convergence to be quite stable across random seeds.
Table 2: Time to evaluate a length $T=10,000$ trained AR GRU using sequential vs parallelized methods. We note the dynamax package [Chang et al., 2023] we used for the parallel Kalman filter implementation in ELK is not optimized for speed, and hence these run times could be further improved.
Algorithm Average time per Newton iteration $\pm$ std (ms) Number of Newton iterations to convergence Total time to convergence (ms)
@@
Similarly, we see that the quasi methods are dramatically faster than their dense counterparts as they are replace $\mathcal{O}(D^{3})$ matrix-matrix multiplication with $\mathcal{O}(D)$ diagonal matrix multiplication.

### v3

@@ Appendix B Experimental Details
B.1 Quasi-DEER for Evaluation
In this section we elaborate on our Experiment 1, discussed in Section 6.1. We closely follow the experimental design in Section 4.1 of Lim et al. [1], including 5 warm-up steps for all timing experiments and a batch size of 16. However, instead of running 5 seeds for 20 repetitions each, we run 20 seeds for 5 repetitions each, to get more coverage over different evaluation (as the timing for each evaluation are observed to be low variance). We also include memory profiling experiments not present in Lim et al. [1]. For these experiments we use 3 random seeds and only one repetition, because the memory usage is extremely stable in between runs. We discus the memory profiling experiments in more detail in Section B.1.3.
For runs with the same specifications (sequence length $T$, hidden state size $D$, and algorithm), we observe that sometimes runs with memory profiling ran out of memory whereas runs with timing profiling did not run out of memory. A difference between our time profiling runs and memory profiling runs was how we handled preallocation of memory. For time profiling, we allowed JAX to preallocate memory because that is how JAX usually runs, and so gives a better indication of wall-clock time in practice. For memory profiling, we did not allow JAX to preallocate memory so we could get a more fine-grained measure of memory usage.
However, JAX provides this following discussion of memory preallocation (see https://jax.readthedocs.io/en/latest/gpu_memory_allocation.html#):
However, [not preallocating memory] is more prone to GPU memory fragmentation, meaning a JAX program that uses most of the available GPU memory may OOM with preallocation disabled.
Because the specifications where the time profile runs stay within memory but the memory runs run out of memory are likely very close to the 16GB threshold, our hypothesis is that this phenomenon is a manifestation of this documented memory fragmentation.
@@ B.1.1 Numerical Precision of DEER and Quasi-DEER
Figure 5: The accuracy of evaluating with parallelized methods (DEER and quasi-DEER) as opposed to sequential evaluation. The parallelized methods converge to the correct trace within numerical precision. The hidden state size is $D=4$ and the sequence length is $T=10,000$.
@@ B.1.2 Different Scaling Regimes Depending on GPU Saturation
In Figure 6, we run the timing benchmarks of Section 6.1 on a wider range of sequence lengths $T$ and hidden state sizes $D$, on a larger GPU (a V100 with 32 GB) and with a smaller batch size of 1. In doing so, we highlight that the parallel nature of DEER and quasi-DEER, as their wall-clock time scales sublinearly in the sequence length $T$ in smaller ($D$, $T$) regimes. However, we note that in the larger regimes considered in our main text and in Lim et al. [1], we often observe linear scaling in the sequence length $T$ for the wall-clock time of DEER and quasi-DEER, even though these algorithms are still faster than sequential evaluation. Figure 6 shows good evidence that these parallel algorithms are suffering from saturation of the GPU, and would benefit from even more optimized parallel hardware
The parallel scan, given sufficiently many processors, scales as $O(\log T)$. As we show in Figure 6, we see this speedup at low model sizes and sequence lengths. Once the processors are saturated, we see a linear increase in the runtime (since the amount of work done is linear), but it is making much more effective use of the GPU, resulting in a constant factor speedup over sequential application at larger model sizes/sequence lengths.
Figure 6: Evaluating an untrained GRU. Sublinear and linear timing regimes for parallelized algorithms. The above experiments were run on a 32 GB V100 with a batch size of 1. As in Figure 2, we use 20 seeds for timing, 3 seeds for memory, and the dashed black line indicates the memory capacity of the GPU (32 GB). We observe that in smaller regimes in $D$ and $T$ that the wall-clock time shows sublinear scaling indicative of the use of parallel algorithms. However, when the GPU becomes saturated, the benefits of parallelization are reduced and we begin to see linear scaling in wall-clock time with $T$.
@@ B.1.3 Memory Profiling Details
B.2 Quasi-DEER for Training
Here we discuss the experimental details for Experiment 2 in Section 6.2. We follow the same experimental setup as in Section 4.3 and Appendix B.3 of Lim et al. [1]. As an aside, we note that the choice of hardware can impact behavior of the algorithms dramatically. For replicability, we run on the same hardware as Lim et al. [1], using a 16GB V100 SXM2. However, we note that if we try to run these same experiments on A100, DEER struggles to converge numerically, although quasi-DEER shows no such difficulty. If we run on a CPU, both DEER and quasi-DEER converge numerically. On balance, on the eigenworms time series classification task, both DEER and quasi-DEER are numerically stable for the most part; the numerical instabilities we have observed for DEER on an A100 are likely specific to some particular detail of JAX/hardware interaction.
In our implementation of quasi-DEER for this experiment, for ease of implementation we do not use the memory efficient version, i.e. we instantiate the full Jacobian and then take the diagonal of it. Nonetheless, we still demonstrate speed gains. Directly implementing the diagonal derivative would likely lead to further speed and memory gains.
The RNN used in this experiment is a 5 layer GRU. When we evaluate this architecture in parallel, we evaluate each layer in parallel using (quasi)-DEER. In Figure 3 (right), we report the number of (quasi)-DEER iterations averaged over all layers and batches. In Figure 7, we report the number of iterations needed for convergence in the last layer only.
Figure 7: The number of Newton iterations needed for the deepest layer of the 5 layer GRU used for the time-series classification task in Section 6.2.
@@ B.3.1 AR GRU Architecture
The architecture is a GRU with hidden states $\mathbf{h}_{t}\in\mathbb{R}^{3}$ and scalar inputs $x_{t}\in\mathbb{R}$. However, at every point in the sequence $t,$ we readout the hidden state $h_{t}\in\mathbb{R}^{3}$ and use it to parameterize a mean $\mu_{t+1}\in\mathbb{R}$ and a variance $\sigma^{2}_{t+1}\in\mathbb{R}_{+}$. We then sample $x_{t+1}$ according to $x_{t+1}\sim\mathcal{N}(\mu_{t+1},\sigma_{t+1}^{2})$; this output $x_{t+1}$ is then fed into as the input to the AR GRU at time step $t+1$ to make the new hidden step $\mathbf{h}_{t+1}$.
This AR GRU is trained using standard sequential evaluation and backpropagation-through-time to produce a noisy sine wave of length 10,000. We train the AR GRU on 1024 traces $\mathbf{x}_{1:T}$ generated from a sine wave with amplitude 10 and white noise applied to each time step, and the training objective is to minimize the the negative log probability of the $\mathbf{x}_{1:T}$.
Once the AR GRU has been trained, it can generate its own trace $\tilde{\mathbf{x}}_{1:T}$ given an initial hidden state $\mathbf{h}_{0}$ and noises $\bm{\epsilon}_{1:T}$.
We note that such a system is Markovian with dimension $D=\dim(\mathbf{h})+\dim(x)$, as together the hidden state $\mathbf{h}_{t}$ and output $x_{t+1}$ determine the next hidden state $\mathbf{h}_{t+1}$ and output $x_{t+2}$. Thus, in the notation of Section 2, a hidden state $\mathbf{s}_{t}$ of the Markovian state space model is $\mathbf{s}_{t}=(x_{t+1},\mathbf{h}_{t})$. Therefore, we can apply fixed-point methods to try to find the correct trace $\mathbf{s}^{*}$ in a parallelized manner instead of autoregressively.
@@ B.3.2 Wall-clock Time Benchmark
The timing experiments were carried out as follows on an Nvidia A100 GPU with 80 GB of GPU memory, using python 3.9 and jax 0.4.11. Exact wall-clock times will vary depending on hardware, software, and implementation.
We ran sequential evaluation of the trained AR GRU to produce noisy sine waves of length $T=10,\!000$, as well as the four parallelized method we consider in this paper (DEER, quasi-DEER, ELK, and quasi-ELK).
We ran 20 different random seeds (which lead to different values of the $\bm{\epsilon}_{1:T}$ and therefore different nonlinear dynamics), and timed each for a total of 4 repetitions (i.e. 80 timing runs per method). We record the wall-clock time needed to evaluate the length $T$ sequence sequentially, as well as wall-clock time, divided by $T$ needed to run $T$ Newton iterations of each of the parallelized methods (thus, we obtain the time per Newton iteration for each of the parallelized methods).
Over these 80 timing runs, the sequential evaluation took an average of 96 milliseconds, with standard deviation of $1.55$ ms. We report the average time per Newton iteration, the total number of iterations needed for convergence, and the total wall-clock time to convergence in Table 2. Note that the third column of Table 2 is the product of the first two columns.
We effectively read the number of Newton iteration to convergence off of the graphs in Figure 4, but find the number of Newton iterations to convergence to be quite stable across random seeds (see Figure 8).
Table 2: Time to evaluate a length $T=10,000$ trained AR GRU using sequential vs parallelized methods. We note the dynamax package [27] we used for the parallel Kalman filter implementation in ELK is not optimized for speed, and hence these run times could be further improved.
Algorithm Time per Newton step (ms, mean $\pm$ std) Newton steps to convergence Total time to convergence (ms)
@@
Similarly, we see that the quasi methods are dramatically faster than their dense counterparts as they are replace $\mathcal{O}(D^{3})$ matrix-matrix multiplication with $\mathcal{O}(D)$ diagonal matrix multiplication. The $\mathcal{O}(D^{3})$ work required by a parallel scan on a dense linear recurrence likely saturates the GPU). We see in Table 2 that individual steps in the dense DEER/ELK are (approximately) a factor of between 3.5 and 30 times slower per step than their quasi (diagonal) variants. However, they take a factor of between 2 and 10 fewer iterations.